# SynSeg: Feature Synergy for Multi-Category Contrastive Learning in Open-Vocabulary Semantic Segmentation

### v1

@@ Introduction
Ablation Study
We examine the effectiveness of the four training objectives by evaluating the impact on segmentation performance of all four loss functions across three representative datasets: Context (Mottaghi et al. 2014), Object (Lin et al. 2014) and Stuff (Caesar, Uijlings, and Ferrari 2018). The results, presented in Tab. 2, show that the combination of all four losses yields the best performance. Removing any single loss leads to a measurable drop in performance, underscoring the complementary contributions of each component in enhancing feature discrimination and segmentation quality. In particular, the inter-category separation loss $L_{sep}$ is the most prominent because the drop is the largest after it is removed.
@@
Visual Effects
In this subsection, we present two types of visual examples to demonstrate the segmentation performance of our method, SynSeg, and also in comparison with existing baselines. The examples are selected from the Context dataset (Mottaghi et al. 2014), as it is a common benchmark supported by all baseline methods.
We begin by presenting a qualitative example of semantic activation maps and the corresponding segmentation masks predicted by our model under multiple threshold settings, as illustrated in Fig. 3. The example involves five semantic categories: bicycle, car, person, grass, and tree. Unlike traditional OVSS where pixel-level predictions are mutually exclusive (Xu et al. 2022; Cha, Mun, and Roh 2023), our semantic activation maps allow a single pixel to be activated by multiple categories simultaneously. This fits the nature in real-world environments, especially when the objects are physically entangled, for example, between the head and the person.
@@
In addition, we conduct visual comparisons between our method and existing methods. Our comparison includes weakly-supervised approaches GroupViT (Xu et al. 2022) and TCL (Cha, Mun, and Roh 2023), and a strong training-free baseline ProxyCLIP (Lan et al. 2024). ProxyCLIP utilizes a large-scale pretrained Vision Foundation Model (VFM) DINO (Jose et al. 2024) for OVSS task. As shown in Fig. 4, our method consistently demonstrates more accurate semantic localization and finer-grained segmentation. The predicted segmentations from our model are better aligned with the object boundaries and exhibit fewer false positives.
@@ Conclusion
We propose a novel approach, SynSeg, for open-vocabulary semantic segmentation, integrating Feature Synergy Structure (FSS) with Multi-Category Contrastive Learning (MCCL). Our approach effectively addresses two key challenges in weakly-supervised OVSS: (1) the lack of explicit inter-category correlation injection during training, and (2) the absence of a suitable way of feature reconstruction tailored for contrastive learning. To address these issues, we introduced MCCL that constructs both intra- and inter-category contrastive learning to enhance the weakly-supervised signal. In addition, FSS generates category-aware features by fusing conditional visual embeddings with semantic-activation maps, enabling finer semantic representation for effective MCCL. Experiments on five benchmarks demonstrate the effectiveness of our method. SynSeg consistently outperforms existing weakly-supervised baselines, achieving state-of-the-art performance.
Technical Appendix
Appendix A Data Pre-Processing
To construct high-quality noun-image pairs from the large-scale CC12M dataset for contrastive learning, we design a two-stage pipeline to extract and filter noun phrases from image captions.
Stage 1: Noun Phrase Extraction
We first extract noun phrases from each caption using the spaCy NLP library (Honnibal and Montani 2017). Given a caption $x$, we extract syntactic noun chunks $\mathcal{N}(x)=\{n_{1},n_{2},\dots,n_{k}\}$ and retain those that contain at least one noun token (with pos_ tag NOUN) to ensure semantic relevance. This step eliminates phrases dominated by modifiers or non-content words.
The result is a mapping from each image to a set of candidate noun phrases:
$\texttt{Image}_{i}\rightarrow\texttt{['dog', 'park', 'south', ...]}$
The extraction process is summarized in Algorithm 1.
Algorithm 1 Noun Phrase Extraction from Captions
Input: Caption file with image-caption pairs
Output: Mapping from image to noun phrases
1:for each line in the caption file do
2:  Parse caption using spaCy
3:  noun_phrases $\leftarrow$ []
4:  for each chunk in parsed.noun_chunks do
5:   Filter tokens with pos_ == NOUN
6:   if at least one noun is retained then
7:    Join tokens and append to noun_phrases
8:   end if
9:  end for
10:end for
11:return image $\rightarrow$ noun_phrases
Stage 2: Generic Phrase Filtering
To enhance the grounding quality of the extracted noun phrases, we further filter out overly generic or non-referential terms such as “reflection”, “southwest” or “background” that lack clear visual correspondence. This is achieved by performing partial string matching against a manually defined exclusion list $\mathcal{E}$. A phrase $p$ is retained only if it contains no substring from $\mathcal{E}$:
$\forall e\in\mathcal{E},\quad e\notin p$
The full filtering procedure is outlined in Algorithm 2.
Algorithm 2 Generic Phrase Filtering
Input: List of noun phrases, exclusion list $\mathcal{E}$
Output: Filtered noun phrase list
1:for each noun phrase $p$ in the list do
2:  if $\forall e\in\mathcal{E},\quad e\notin p$ then
3:   Keep $p$
4:  else
5:   Discard $p$
6:  end if
7:end for
8:return filtered noun phrases
The final output is a cleaned set of noun-image pairs that serve as anchors for downstream contrastive training.
Appendix B Experimental Setup
Experimental Enviroment
[…]
The Python environment is based on Python 3.9.20. Major software libraries and dependencies include:
  • PyTorch 1.12.1 + CUDA 11.6
  • TorchVision 0.13.1 + CUDA 11.6
  • MMCV 2.0.0rc3
  • MMSegmentation 1.0.0rc4
  • SpaCy 3.8.0
To ensure reproducibility, we fix the random seed to 42 in all experiments. And each experiment (including training and evaluation) was run independently on a single RTX 4090 GPU, as our model and training recipes are sufficiently lightweight. All reported results are averaged over three independent runs.
Hyperparameter Configuration
The hyperparameter settings for the main reported experiments are as follows:
  • Loss weights: $\lambda_{1}=1$, $\lambda_{2}=1$, $\lambda_{3}=10$, $\lambda_{4}=1$
  • Learning rate: $1\times 10^{-5}$
  • Weight decay: $1\times 10^{-4}$
  • Batch size: 128
  • Number of epochs: 1
  • Optimizer: SGD
These values are used consistently unless otherwise specified.
Appendix C Dataset Details
We evaluate our method on five widely used open-vocabulary semantic segmentation benchmarks: PASCAL VOC (Everingham et al. 2010), Pascal Context (Mottaghi et al. 2014), Cityscapes (Cordts et al. 2016), COCO Object, and COCO Stuff (Lin et al. 2014; Caesar, Uijlings, and Ferrari 2018). These datasets collectively cover a diverse range of semantic granularity and scene complexity, which makes them suitable for comprehensive evaluation under open vocabuary conditions.
PASCAL VOC (VOC).
PASCAL VOC is a classical image-based dataset for classification, object detection and segmentation. We use the latest PASCAL VOC 2012 version, which contains 21 semantic categories with an average of 2.5 labeled instances per image. Following prior work, unlabeled regions are treated as a separate background class during evaluation.
Pascal Context (Context).
Pascal Context expands VOC annotations to include 59 object and stuff categories with denser labeling. On average, each image contains 4.8 annotated categories, offering more complex scene understanding challenges. Only labeled pixels are used for evaluation, and background is not treated as an explicit class.
Cityscapes (City).
Cityscapes focuses on urban street scenes, containing 19 finely annotated semantic categories. The validation set of Cityscapes features high-resolution images with an average of 12.0 labeled categories per image. As with Context, evaluation is performed only on annotated regions. Its strong domain specificity makes it useful for testing model robustness under structured real-world conditions.
COCO Object (Object).
COCO is a famous dataset built in 2014 and updated in 2017, which concentrates on Commen Objects in Context. COCO Object includes 80 object classes from the COCO dataset, annotated with instance-level masks. The evaluation treats unlabeled areas as an extra background class. The average number of labeled targets per image is 7.2, making it moderately challenging in both category size and visual diversity.
COCO Stuff (Stuff).
COCO Stuff extends COCO annotations with 91 stuff and 80 object categories (171 total), densely covering nearly every pixel in an image. The validation set has an average of 8.4 semantic regions per image. Evaluation focuses only on labeled pixels.
Summary .
These datasets collectively evaluate segmentation performance from different semantic settings with wide recognition. Table 3 summarizes their statistics.
Dataset #Categories Labels/Image Background
PASCAL VOC 21 2.5 Yes
Pascal Context 59 4.8 No
Cityscapes 19 12.0 No
COCO Object 81 7.2 Yes
COCO Stuff 171 8.4 No
Table 3: Summary of dataset statistics and evaluation settings.

### v2

@@ 5.2 Main Results
5.3 Ablation Study
We examine the effectiveness of the four training objectives by evaluating the impact on segmentation performance of all four loss functions across three representative datasets: Context [31], Object [26] and Stuff [1]. The results, presented in Tab. 3, show that the combination of all four losses yields the best performance. Removing any single loss leads to a measurable drop in performance, underscoring the complementary contributions of each component in enhancing feature discrimination and segmentation quality. In particular, the inter-category separation loss $L_{sep}$ is the most prominent because the drop is the largest after it is removed.
@@ 5.3 Ablation Study
5.4 Visual Effects
In this subsection, we present two types of visual examples to demonstrate the segmentation performance of our method, SynSeg, and also in comparison with existing baselines. The examples are selected from the Context dataset [31], as it is a common benchmark supported by all baseline methods.
We begin by presenting a qualitative example of semantic activation maps and the corresponding segmentation masks predicted by our model under multiple threshold settings, as illustrated in Fig. 3. The example involves five semantic categories: bicycle, car, person, grass, and tree. Unlike traditional OVSS where pixel-level predictions are mutually exclusive [50, 4], our semantic activation maps allow a single pixel to be activated by multiple categories simultaneously. This fits the nature in real-world environments, especially when the objects are physically entangled, for example, between the head and the person.
@@ 5.4 Visual Effects
In addition, we conduct visual comparisons between our method and existing methods. Our comparison includes weakly-supervised approaches GroupViT [50] and TCL [4], and a strong training-free baseline ProxyCLIP [23]. ProxyCLIP utilizes a large-scale pretrained Vision Foundation Model (VFM) DINO [17] for OVSS task. As shown in Fig. 4, our method consistently demonstrates more accurate semantic localization and finer-grained segmentation. The predicted segmentations from our model are better aligned with the object boundaries and exhibit fewer false positives.
@@ 6 Conclusion
We propose a novel approach, SynSeg, for end-to-end open-vocabulary semantic segmentation, integrating Feature Synergy Structure (FSS) with Multi-Category Contrastive Learning (MCCL). Our approach effectively addresses two key challenges in weakly-supervised OVSS: (1) the lack of explicit inter-category correlation injection during training, and (2) the absence of a suitable way of feature reconstruction tailored for contrastive learning. To address these issues, we introduced MCCL that constructs both intra- and inter-category contrastive learning to enhance the weakly-supervised signal. In addition, FSS generates category-aware features by fusing conditional visual embeddings with semantic-activation maps, enabling finer semantic representation for effective MCCL. Experiments on five benchmarks demonstrate the effectiveness of our method. SynSeg consistently outperforms existing weakly-supervised baselines in OVSS, achieving state-of-the-art performance.